OpenStudy (anonymous):

what are critical points where f prime is 0?? on x^4/5(x+3)

OpenStudy (helder_edwin):

u want to find the critical points of \[ \large f(x)=\frac{x^4}{5(x+3)} \] ?

OpenStudy (anonymous):

hello? ru there?

OpenStudy (anonymous):

critical point is defined as where f' = 0 or f' does not exist

OpenStudy (anonymous):

within the domain of f, of course....

OpenStudy (anonymous):

noooo x^5/4

OpenStudy (anonymous):

can you draw it?

OpenStudy (anonymous):

noooo, i'm failing to

OpenStudy (anonymous):

i mean use the "draw" button to show the function.....

OpenStudy (anonymous):

x^4/5 * (x+3)

OpenStudy (anonymous):

\[\huge f(x)=x^{\frac{4}{5}}(x+3) \] like that?

OpenStudy (anonymous):

x^4/5 * (x+3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there is the graph! open the link

OpenStudy (anonymous):

i was not asking for the graph... i was asking that use the draw botton to write the function so that it would be clear what function we're talking about...

OpenStudy (anonymous):

but we already have it , you confirmed it....

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

best way to take the derivative here is to just multiply it out and use power rule, sum rule for derivatives... \[\huge f(x)=x^{\frac{4}{5}}(x+3)=x^{\frac{4}{5}}\cdot x+x^{\frac{4}{5}}\cdot 3 \] \[\huge f(x)=x^{\frac{9}{5}}+3x^{\frac{4}{5}} \] can you take the derivative from here?

OpenStudy (anonymous):

i got that part right... do i derive that?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

what u got for the derivative?

OpenStudy (anonymous):

mmm

OpenStudy (anonymous):

9/5x^(4/5) + 12/5x^9(

OpenStudy (anonymous):

+ 12/5x^9(

OpenStudy (anonymous):

its freezing

OpenStudy (anonymous):

the last part is 12/5x^(-1/5)

OpenStudy (anonymous):

that last part should be + (12/5)x^(-1/5) , right?

OpenStudy (anonymous):

ok... good...

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\huge f(x)=x^{\frac{4}{5}}(x+3)=x^{\frac{9}{5}}+3x^{\frac{4}{5}} \] \[\huge f'(x)=\frac{9}{5}x^{\frac{4}{5}}+\frac{12}{5}x^{-\frac{1}{5}} \] \[\huge f'(x)=\frac{9}{5}x^{\frac{4}{5}}+\frac{12}{5}x^{-\frac{1}{5}} = \frac{3(3x+4)}{5x^{\frac{1}{5}}}\] ok... i just simplified your derivative... set that equal to zero, what u got?

OpenStudy (anonymous):

oh, btw, the graph you attached is NOT the graph of the function we're working on.... \(\large f(x) = x^{4/5}(x+3) \)